OpenStudy (anonymous):

If u = 8 + 4i and v = 3i + 6, then what is u – v ? A)2 + 2i B)5– 2i C)1 + i D)2 + i E)2 – i

OpenStudy (anonymous):

u = 8 + 4i v = 3i + 6 u-v = (8 + 4i) - (3i + 6) = 8 - 6 + 4i - 3i = i + 2 So, your answer is D.

OpenStudy (anonymous):